Solve the following system of equations for x and y
3x+5y=21
2x-y=1

Answer by rfer(16322)

3x+5y=21
2x-y=1, y=2x-1
--------------------
3x+5(2x-1)=21
3x+10x-5=21
13x=26
x=2
-----------------
2(2)-y=1
4-y=1
-y=-3
y=3
--------------------
proof
3(2)+5(3)=21
6+15=21